# POp-GS: Next Best View in 3D-Gaussian Splatting with P-Optimality

### v1

@@ Abstract
In this paper, we present a novel algorithm for quantifying uncertainty and information gained within 3D Gaussian Splatting (3D-GS) through P-Optimality. While 3D-GS has proven to be a useful world model with high-quality rasterizations, it does not natively quantify uncertainty. Quantifying uncertainty in parameters of 3D-GS is necessary to understand the information gained from acquiring new images as in active perception, or identify redundant images which can be removed from memory due to resource constraints in online 3D-GS SLAM. We propose to quantify uncertainty and information gain in 3D-GS by reformulating the problem through the lens of optimal experimental design, which is a classical solution to measuring information gain. By restructuring information quantification of 3D-GS through optimal experimental design, we arrive at multiple solutions, of which T-Optimality and D-Optimality perform the best quantitatively and qualitatively as measured on two popular datasets. Additionally, we propose a block diagonal approximation of the 3D-GS uncertainty, which provides a measure of correlation for computing more accurate information gain, at the expense of a greater computation cost.
@@
Figure 1: In this paper, we propose a novel method for calculating information gain from image in 3D Gaussian Splatting. In the above images, each method is provided a set of one hundred candidate views on scenes from the Blender dataset, and selects ten views to train a 3D-GS model on. Compared to state-of-the-art, our method more accurately estimate the information value of images to a trained 3D-GS model. In this figure we demonstrate block D-Optimality, however our derivation also provides multiple solutions discussed later.
@@ 1 Introduction
In order to operate in a novel environment, robots must be capable of quantifying uncertainty in their surroundings due to occlusions and unobserved regions, as well as quantifying information gained by exploring new areas. While many prior methods for map representations were built on a metric space with natural uncertainty quantification, such representations result in discretization errors of the environment [36]. Recently, 3D-Gaussian Splatting (3D-GS) was proposed as a method for high-quality novel-view synthesis [14], which captivated the attention of the robotics community due to its explicit world model representation, with the potential to function as a map [6, 23, 40].
@@
Given a set of views of a scene, 3D-GS learns to render novel views from any angle through gradient descent. 3D ellipsoids are iteratively fine-tuned, which can then be rasterized to a 2D image through a process known as Gaussian Splatting. Several works have shown that the 3D ellipsoids can be extended to include features beyond just color, such as the category of objects the ellipsoid represents, resulting in an improved level of semantic scene understanding [39, 30]. However, 3D-GS does not natively quantify uncertainty, which leads to issues when determining whether an image has been seen before, or when quantifying the information gained by perceiving a new view. While several works have studied this problem for neural radiance fields (NeRF) [8, 26, 25], the explicit representation of 3D-GS presents a unique challenge.
@@
Since the conception of 3D-GS, several works have sought to quantify uncertainty directly upon a trained 3D-GS model [9, 13, 18]. Of these works, one solution which has appeared effective is relating the per-pixel gradient of the 3D-GS parameters to the information or contribution of the parameters. Notably, FisherRF [13] derived a solution for information quantification of images in 3D-GS through Fisher Information [16], which approximated uncertainty through the diagonal elements of an information matrix. FisherRF has since been applied to mobile manipulation [33] and active perception [21], however it does not leverage prior literature in effective functions for active perception [15, 32], and does not consider any correlation between parameters.
@@
Classically, the problem of quantifying information gain from views is known as active perception in robotics literature, and has been studied extensive in probabilistic robotics applications [22, 7]. More generally, active perception can be solved through application of optimal experimental design [19], which defines solutions for identifying the most informative design parameters for an experiment in order to learn about a set of unknown parameters [20, 10, 11, 17]. Although optimal experimental design research provides insight towards how the information value of an image may be quantified in 3D-GS, application of P-Optimality metrics ultimately relies on a covariance matrix which is not directly available in 3D-GS [15].
@@ 2.2 Active Perception
Active perception is a well-studied problem in robotics literature which seeks to identify the optimal path to improve a map. Since capturing new views in the real world requires robot traversal, a significant amount of research has focused on optimal solutions to determining which view-points are most valuable. One successful approach is the application of P-optimality [15], which defines a class of optimal experimental design solutions based on the value of an exponential. By formulating the information gain problem in terms of the eigenvalues of an information matrix, P-optimality provides methods for computing the information value of measurements.
Early research in optimal decision-taking for Simultaneous Localization and Mapping problems used T-optimality due to efficient computation as the trace of the covariance matrix [32, 24], eliminating the need to compute the eigenvalues. On the other hand, recent research has focused on D-optimality as a reliable metric for an Optimal Experimental Design [29, 4, 27, 28]. As advocated in [3], D-optimality seems to be a more rewarding metric towards the fulfilling of an active mapping session. The recent success of D-optimality can also be explained by its monotonicity property in active mapping scenarios [31], which guarantees that uncertainty increases monotonically as a robot move through the scene. In this work, we propose to expand upon the formulation from FisherRF to incorporate parameter correlation and develop a more general solution which allows for application of classical optimal experimental design techniques.
@@ 3.2 Information Gain through Optimal Experimental Design
Figure 2: The Hessian matrix captures the information content of each parameter in the trained 3D-GS model, approximated as per-pixel gradients over the training set of images. Since a trained 3D-GS model may contain millions of parameters, we approximate the Hessian matrix through a main diagonal and block diagonal approximation.
Assuming that the optimal solution $\bm{\theta}_{*}$ is unbiased, then it follows that $\mathbb{E}[\bm{e}_{*}]=\bm{0}$ and $\mathbb{E}[\bm{e}_{*}\bm{e}_{*}^{T}]=\bm{\sigma}_{e}\cdot\mathbb{I}$, leading to $\mathbb{E}[\Delta\bm{\theta}]=\bm{0}$ and:
@@
Equivalent Formulae $\frac{1}{l}\text{tr}\begin{pmatrix}\bm{\Sigma}_{i}\end{pmatrix}=\frac{1}{l}\sum_{k}^{l}\lambda_{k}$ $\frac{1}{l}\text{tr}\begin{pmatrix}\bm{H}_{i}\end{pmatrix}=\frac{1}{l}\sum_{k}^{l}\lambda_{k}^{-1}$ |$n]{|\bm{\Sigma}_{\theta}|}=\exp\begin{pmatrix}\frac{1}{l}\sum_{k}^{l}\log\lambda_{k}\end{pmatrix}$ $\min_{\lambda_{k}}$ $\max_{\lambda_{k}}$
Meaning Average Variance Average Information Volume of covariance hyper-ellipsoid Single extreme eigenvalue
@@ 3.2.1 Uncertainty Decrease due to an Added Image
Given that we have an inverse relation between the covariance matrix and the Hessian, we need to find the optimal solution to:
$\min_{i}U\begin{pmatrix}\bm{H}_{i}^{-1}\end{pmatrix}=\min_{i}U\begin{pmatrix}\bm{\Sigma}_{i}\end{pmatrix},$ (12)
where $U\begin{pmatrix}\cdot\end{pmatrix}$ is a measure of uncertainty for the covariance matrix. There are multiple ways we can define the uncertainty function. In this work, we rely on the Theory of Optimal Experimental Design [29, 15], which defines the P-Optimality uncertainty metric as:
$U_{p}(\bm{\Sigma}_{i})=\begin{pmatrix}\frac{1}{l}\text{trace}\begin{pmatrix}\bm{\Sigma}_{i}\end{pmatrix}\end{pmatrix}^{\frac{1}{p}},$ (13)
where $p$ is an integer. Depending on the chosen value for $p$, the uncertainty function can have some special properties [27], as detailed in Table 1. Note that matrix inversion can be avoided by inverting the eigenvalues instead of the information matrix, resulting in the same value.
@@ 3.3 Approximating the Covariance
Therefore, in order to calculate P-Optimality we propose to first calculate the Hessian matrix and next calculate P-Optimality functions over the eigenvalues of the Hessian matrix. However, computation of the eigenvalues has a computational complexity of $\mathcal{O}(n^{3})$ and memory complexity of $\mathcal{O}(n^{2})$, which is intractable for 3D-GS models where the number of parameters $n$ may be in the millions. Therefore, we propose two approximations of the Hessian matrix to save memory and computation.
Simple Diagonal: First, based off the work of FisherRF, we propose a simple approximation of the Hessian matrix which only incorporates the main diagonal. FisherRF derives the approximation from a Laplace approximation [5], where the covariance matrix is approximated as the main diagonal plus a small regularizing constant $\lambda_{\theta}$:
$\Sigma_{\theta}\approx\text{diag}(\Sigma_{\theta})+\lambda_{\theta},$ (14)
@@
Block Diagonal: In order to capture some of the correlation between 3D-GS parameters, we propose to approximate the Hessian matrix as a block diagonal matrix where each block diagonal element contains the parameters of a single ellipsoid. Please note that a block diagonal approximation has also been proposed within 3D-GS, however the approximation was applied to the task of pruning 3D-GS models to remove redundant ellipsoids [9]. Our insight is that the parameters of the same ellipsoid are most likely to be correlated, and a block diagonal matrix can be processed in parallel on a GPU for efficient computation of information gain.
@@ 3.4 Batch Selection
Sometimes, the information content of multiple images may be valuable to measure at a time. For example, a 3D-GS model may be provided hundreds of images as training input, which may contain redundant images. Batch image selection can be used to measure the information content of images and select a batch of the most important images. In this problem setup, the goal is to identify the information value of a set of images. Following FisherRF, we implement a simple approach which iteratively adds or removes a candidate image, updates the Hessian, and repeats the process without additional training. While simple, this batch information implementation does capture the redundancy of views, as the change in the Hessian is reflected when a candidate image is added.
@@ 4 Results
In this section, we study the effectiveness of our method on quantifying information gained from obtaining new images. Following the experimental setup of FisherRF, we quantitatively and qualitatively evaluate our method against baselines on the task of single view selection, where a 3D-GS model is trained by iteratively selecting the most informative proposal view and fitting the model to a set of candidate view. Next, we compare our method against the same baselines on a similar experimental setup, where multiple images are chosen at a time for a batch view selection problem. Third, we compare the ability of each method to quantify uncertainty in novel views by studying the correlation of information gain with reconstruction metrics. Last, we perform ablation studies on the parameters of the 3D-GS model to identify the most important parameters for calculating information gain.
Baselines: We compare against the recently published FisherRF, which calculates the next best view from the trace of the proposal image. While FisherRF compared against a random and ActiveNeRF baseline [26], we omit these baselines as they performed worse than FisherRF in their experiments, and instead focus on comparing our results with FisherRF. We also add a uniform sampling baseline, which is slightly different from the implementation of random in FisherRF that we found was incorrectly implemented. Instead, our uniform sampling baseline samples views uniformly from the training set, which results in a high coverage of the views in the test set, especially if the training and testing views are similarly distributed. As one of the merits of our approach is a more general solution with optimal experimental design, we implement A, D, E, and T Optimality baselines for comparison. All baselines are compared over the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [35], and LPIPS metrics [38].
@@ 4.1 Single View Selection
Twenty Views: The twenty view experimental setup follows a very similar approach, of iteratively adding views and training for $100v$ iterations before selecting the next view. However, this setup begins with $v=4$ training views, adds images until $v=20$, and trains until a cumulative total of $21,000$ training steps.
Experimental results on the MIP dataset can be found in Table 4. Similar to the ten view experiment, we find that uniform sampling slightly outperforms FisherRF while T and D Optimality achieve the highest performance. Additionally, block diagonal D-Optimality improves the structural quality of the renderings due to improved information quantification.
@@ 4.2 Batch View Selection
Iterative: In the first experiment on batch view selection we follow the experimental set-up of FisherRF. 3D-GS models are trained on the same set of $4$ initial views, and each method iteratively trains and selects new view to add. However, in this experiment models are trained for $150v$ iterations and select 4 views at a time. 3D-GS models are trained until a cumulative total of $10,000$ training steps for comparison. Results on the Mip-Nerf360 dataset are shown in Table 5, and similar to previous experiments demonstrate superior performance of D and T optimality. Additionally, the block diagonal approximation improves structural quality measured by SSIM and LPIPS metrics. Note that due to the iterative view selection and training, this experiment may not reward batch view diversity, motivating our next experiment.
Table 6: Results on View Compression on Blender Dataset.
@@
Compression: To further study batch view selection, we train a 3D-GS model with all views then select the ten most informative views on the trained model with each method. The selected views are then used to train a 3D-GS model, which is evaluated and compared as a measure of batch information quantification. Results on the Blender dataset are summarized in Table 6, demonstrating low performance of FisherRF, A Optimality and E Optimality which select similar views. Instead, T and D Optimality select informative and different views resulting in a large performance gap.
4.3 Uncertainty Correlation
Since the goal of information quantification is to identify images which would contribute the most information to our trained 3D-GS model, we expect that the 3D-GS model would create poorly rendered images at the most informative candidate views. Therefore, as another test of our proposed method of quantifying information gain, we create a sparsification plot [12] which identifies the correlation between image information gain and image render quality shown in Fig. 5.
The sparsification plot is created by first training a 3D-GS model on ten randomly selected images for $2,000$ iterations. Each method, using the same random seed and trained model, sorts test views by expected information gain. Views are separated into ten bins, where the first bin contains the most informative views, expected to have the lowest quality indicated by a low PSNR. Iteratively, bins are added to the evaluation set which should result in increased average render quality.

### v2

@@ Abstract
In this paper, we present a novel algorithm for quantifying uncertainty and information gained within 3D Gaussian Splatting (3D-GS) through P-Optimality. While 3D-GS has proven to be a useful world model with high-quality rasterizations, it does not natively quantify uncertainty or information, posing a challenge for real-world applications such as 3D-GS SLAM. We propose to quantify information gain in 3D-GS by reformulating the problem through the lens of optimal experimental design, which is a classical solution widely used in literature. By restructuring information quantification of 3D-GS through optimal experimental design, we arrive at multiple solutions, of which T-Optimality and D-Optimality perform the best quantitatively and qualitatively as measured on two popular datasets. Additionally, we propose a block diagonal covariance approximation which provides a measure of correlation at the expense of a greater computation cost.
@@
Figure 1: We propose a novel method for calculating information gain from image in 3D Gaussian Splatting. In the above images, each method is provided a set of one hundred candidate views on scenes from the Blender dataset, and selects ten views to train a 3D-GS model on. Compared to state-of-the-art, our method more accurately estimate the information value of images to train a 3D-GS model. In this figure we demonstrate block D-Optimality, however our derivation also provides multiple solutions discussed later.
@@ 1 Introduction
In order to operate in a novel environment, robots must be capable of quantifying uncertainty in their surroundings due to occlusions and unobserved regions, as well as quantifying information gained by exploring new areas. While many prior methods for map representations were built on a metric space with natural uncertainty quantification, such representations result in discretization errors of the environment [39]. Recently, 3D-Gaussian Splatting (3D-GS) was proposed as a method for high-quality novel-view synthesis [15], which captivated the attention of the robotics community due to its explicit world model representation, with the potential to function as a map [7, 24, 43, 38].
@@
Given a set of views of a scene, 3D-GS learns to render novel views from any angle through gradient descent. 3D ellipsoids are iteratively fine-tuned, which can then be rasterized to a 2D image at novel views. Several works have shown that the 3D ellipsoids can be extended to include features beyond just color, such as the category of objects the ellipsoid represents, resulting in an improved level of semantic scene understanding [42, 31]. However, 3D-GS does not natively quantify uncertainty, which leads to issues when determining whether an image has been seen before, or when quantifying the information gained by perceiving a new view. While several works have studied this problem for neural radiance fields (NeRF) [9, 27, 26], the explicit representation of 3D-GS presents a unique challenge.
@@
Since the conception of 3D-GS, several works have sought to quantify uncertainty directly upon a trained 3D-GS model [10, 14, 19]. Of these works, one solution which has appeared effective is relating the per-pixel gradient of the 3D-GS parameters to the information or contribution of the parameters. Notably, FisherRF [14] derived a solution for information quantification of views in 3D-GS through a diagonal approximation of Fisher Information [17]. FisherRF has since been applied to mobile manipulation [34] and active perception [22], however it does not leverage prior literature in effective functions for active perception [16, 33] or consider any correlation between parameters.
@@
Classically, the problem of quantifying information gain from views is known as active perception in robotics literature, and has been studied extensive in probabilistic robotics applications [23, 8]. More generally, active perception can be solved through application of optimal experimental design [20], which defines solutions for identifying the most informative design parameters for an experiment in order to learn about a set of unknown parameters [21, 11, 12, 18]. One particular solution to experimental design is the use of P-Optimality, which specifies a class of solutions depending on the covariance matrix and the choice of P. P-Optimality has been successfully applied in robotics to keyframe selection, optimal loop closure, and active perception, however has not been explored in 3D-GS [30, 29, 3].
@@ 2.2 Active Perception
Active perception is a well-studied problem in robotics literature which seeks to identify the optimal path to improve a map. Since capturing new views in the real world requires robot traversal, a significant amount of research has focused on optimal solutions to determining which view-points are most valuable. One successful approach is the application of P-optimality [16], which defines a class of optimal experimental design solutions based on functionals of the covariance matrix which vary with the choice of an integer $p$. P-Optimality (P-Opt.) has been widely and successfully applied to SLAM prior to the conception of 3D-GS in keyframe selection, optimal loop closure, and active perception [30, 29, 3].
Early research in optimal decision-taking for Simultaneous Localization and Mapping problems used T-optimality due to efficient computation as the trace of the covariance matrix [33, 25], eliminating the need to compute the eigenvalues. On the other hand, recent research has focused on D-optimality as a reliable metric for Optimal Experimental Design [30, 4, 28, 29], especially in active mapping [3]. The recent success of D-optimality can be explained by its monotonicity property in active mapping scenarios [32], which guarantees that uncertainty increases monotonically as a robot moves through the scene. Additionally, from an information theoretic perspective, differential entropy of a multivariate Gaussian is proportional to the determinant of the covariance matrix, which is captured by D-optimality [5]. In this work, we propose to expand upon the formulation from FisherRF to incorporate parameter correlation and develop a more general solution which allows for application of classical optimal experimental design techniques.
@@ 3.2 Information Gain through Optimal Experimental Design
Figure 2: The Hessian matrix captures the information content of each parameter in the trained 3D-GS model, approximated as per-pixel gradients over the training set of images. Since a trained 3D-GS model may contain millions of parameters, we approximate the Hessian matrix through a block or main diagonal.
Assuming that the optimal solution $\bm{\theta}_{*}$ is unbiased, then it follows that the expected residual $\mathbb{E}[\bm{e}_{*}]=\bm{0}$ and $\mathbb{E}[\bm{e}_{*}\bm{e}_{*}^{T}]=\bm{\sigma}_{e}\cdot\mathbb{I}$, leading to $\mathbb{E}[\Delta\bm{\theta}]=\bm{0}$ and:
@@
Equivalent Formulae $\frac{1}{l}\text{tr}\begin{pmatrix}\bm{\Sigma}_{i}\end{pmatrix}=\frac{1}{l}\sum_{k}^{l}\lambda_{k}$ $\left(\frac{1}{l}\text{tr}\begin{pmatrix}\bm{H}_{i}\end{pmatrix}\right)^{-1}=\left(\frac{1}{l}\sum_{k}^{l}\lambda_{k}^{-1}\right)^{-1}$ |$l]{|\bm{\Sigma}_{i}|}=\exp\begin{pmatrix}\frac{1}{l}\sum_{k}^{l}\log\lambda_{k}\end{pmatrix}$ $\min_{\lambda_{k}}$ $\max_{\lambda_{k}}$
Meaning Average Variance Harmonic Mean Variance Volume of covariance hyper-ellipsoid Single extreme eigenvalue
@@ 3.2.1 Uncertainty Decrease due to an Added Image
For exponential likelihoods, there is an asymptotic inverse relationship in the maximum likelihood estimator between the Hessian and covariance matrices [36]. Therefore, our goal is to maximize the information $\mathcal{I}\begin{pmatrix}\cdot\end{pmatrix}$ obtained from the i-th image, which can be found by minimizing the uncertainty $U\begin{pmatrix}\cdot\end{pmatrix}$:
$\text{arg}\max_{i}\mathcal{I}\begin{pmatrix}\bm{H}_{i}\end{pmatrix}$ $=\text{arg}\min_{i}U\begin{pmatrix}\bm{\Sigma}_{i}\end{pmatrix}$
$=\text{arg}\min_{i}U\begin{pmatrix}\bm{H}_{i}^{-1}\end{pmatrix}.$ (12)
In this work, we rely on the Theory of Optimal Experimental Design [30, 16], which defines the P-Optimality uncertainty metric as:
$U_{p}(\bm{\Sigma}_{i})=\begin{pmatrix}\frac{1}{l}\text{trace}\begin{pmatrix}\bm{\Sigma}_{i}^{p}\end{pmatrix}\end{pmatrix}^{\frac{1}{p}},$ (13)
where $p$ is an integer. Depending on the chosen value for $p$, the uncertainty function can have some special properties [28], as detailed in Table 1.
@@ 3.3 Approximating the Covariance
In practice fitted 3D-GS models may contain millions of parameters, which is intractable due to the cubic computational and quadratic memory complexity of computing the eigenvalues. Therefore, we propose two approximations of the Hessian matrix to save memory and computation.
Simple Diagonal: First, following the work of FisherRF we propose a simple diagonal approximation of the covariance matrix. FisherRF derives the approximation from a Laplace approximation [6], where the covariance matrix is approximated as the main diagonal plus a small regularizing constant $\lambda_{\theta}$:
$\Sigma_{\theta}\approx\text{diag}(\Sigma_{i})+\lambda_{\theta},$ (14)
@@
Block Diagonal: In order to capture some of the correlation between 3D-GS parameters, we propose to approximate the Hessian matrix as a block diagonal matrix where each block diagonal element contains the parameters of a single ellipsoid. Please note that a block diagonal approximation has also been proposed within 3D-GS, however the approximation was applied to the task of pruning 3D-GS models to remove redundant ellipsoids [10]. Our insight is that the parameters of the same ellipsoid are most likely to be correlated, and a block diagonal matrix can be processed in parallel on a GPU for efficient computation.
@@ 3.4 Batch Selection
In practical applications it may be valuable to measure information gain over a set of candidate images, such as along a trajectory or in keyframe selection. Following FisherRF, we implement a simple approach which iteratively adds or removes the most optimal candidate image, updates the Hessian, and repeats the process without additional training. While simple, this batch information implementation does capture the redundancy of views, as the change in the Hessian is reflected when a candidate image is added.
@@ 4 Results
In this section, we study the effectiveness of our method on quantifying information gained from obtaining new images. Following the experimental setup of FisherRF, we quantitatively and qualitatively evaluate our method against baselines on the task of single view selection, where a 3D-GS model is trained by iteratively selecting the most informative proposal view and fitting the model to a set of candidate view. Next, we compare our method against the same baselines on batch view selection. Third, we compare the ability of each method to quantify uncertainty in novel views by studying the correlation of information gain with reconstruction metrics. Last, we perform ablation studies on the parameters of the 3D-GS model to identify the most important parameters for calculating information gain.
Baselines: We compare against the recently published FisherRF [14], which calculates information gain as:
$\mathcal{I}(\bm{H}_{i})=\text{tr}\begin{pmatrix}(\bm{J}_{i}^{T}\bm{J}_{i})\bm{H}^{-1}_{-}\end{pmatrix}$ (15)
with matrices modeled by the simple diagonal approximation. While FisherRF compared against a random and ActiveNeRF baseline [27], we omit these baselines as they performed worse than FisherRF in their experiments, and instead focus on comparing our results with FisherRF. We also add a uniform sampling baseline, which is slightly different from the implementation of random in FisherRF that we found was incorrectly implemented. Instead, our uniform sampling baseline samples views uniformly from the training set, which results in a high coverage of the views in the test set, especially if the training and testing views are similarly distributed. As one of the merits of our approach is a more general solution with optimal experimental design, we implement A, D, E, and T Optimality baselines for comparison. All baselines are compared over the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [37], and LPIPS metrics [41].
@@ 4.1 Single View Selection
Twenty Views: The twenty view experimental setup follows a very similar approach, of iteratively adding views and training for $100v$ iterations before selecting the next view. However, this setup begins with $v=4$ training views, adds images until $v=20$, and trains until a cumulative total of $21,000$ training steps. Experimental results on the MIP dataset can be found in Table 4. Similar to the ten view experiment, we find that uniform sampling slightly outperforms FisherRF while T and D Optimality achieve the highest performance with an improvement from the block diagonal approximation.
@@ 4.2 Batch View Selection
Iterative: In the first experiment on batch view selection we follow the experimental set-up of FisherRF. The procedure is similar to single view selection, however models begin with $4$ training views, are trained for $150v$ iterations between view selections, and select 4 views at a time until $20$ views are obtained. All models are trained for a cumulative total of $10,000$ training steps. Results on the Mip-Nerf360 dataset are shown in Table 5, and similar to previous experiments demonstrate superior performance of D and T optimality. Additionally, the block diagonal approximation improves structural quality measured by SSIM and LPIPS metrics. Note that due to the iterative view selection and training, this experiment may not reward batch view diversity, motivating our next experiment.
Table 6: Results on Keyframe Selection on Blender Dataset.
@@
Keyframe Selection: To further study batch view selection in a setting more similar to SLAM applications, we compare each approach on keyframe selection. All methods are provided the same pre-trained 3D-GS model and select ten keyframes without replacement from a set of views. The selected views are then used to re-train a new 3D-GS model, which is evaluated and compared as a measure of batch information quantification. Results on the Blender dataset are summarized in Table 6, demonstrating low performance of FisherRF, A Optimality and E Optimality which select similar views. Instead, T and D Optimality select informative and different views resulting in a large performance gap.
4.3 Correlation with Render Quality
Intuitively, we expect information gain and rasterization quality at view points to be inversely related. For instance, if a 3D-GS model has only been trained on the front side of a chair, view points from the back side would have poor renderings while providing high information gain to the 3D-GS model. Therefore, as another test of our proposed method of quantifying information gain, we create a sparsification plot [13] to study the inverse relationship between image information gain and image render quality.
The sparsification plot in Fig. 5 is created by first training a 3D-GS model on ten randomly selected images for $2,000$ iterations. Each method, using the same random seed and trained model, sorts candidate views by expected information gain. At decile increments, the cumulative average reconstruction quality of views is calculated and plotted for each method. Reading from left to right, the plot indicates the average reconstruction quality of the most informative views. Due to the inverse nature between image uncertainty and information gain, we would expect the most informative candidate views (left) to have low reconstruction quality.